# D-LEAF: Localizing and Correcting Hallucinations in Multimodal LLMs via Layer–to-head Attention Diagnostics

### v1

@@ C.1 LogitLens
C.2 Wilcoxon Signed-rank Test
@@ C.3 Isotonic Regression
C.4 Spearman Correlation Coefficient
@@ D.1 Empirical Analysis on Entropy and Focus
We randomly sampled 500 images from the COCO2014 validation set and, for each image, extracted the model’s attention matrices when generating hallucinated versus ground‑truth tokens, respectively. We then computed the attention scores over the image region for both cases. The results are shown in Figure 8. We observe that, in both models, across all layers, image‑region attention for ground‑truth tokens is higher than hallucinated tokens.
@@
Similarly, after applying softmax normalization to the attention matrices for ground‑truth and hallucinated tokens, we computed the mean image‑region entropy across all attention heads at each layer and plotted their per‑layer difference as a bar chart in Figure 9. Because the raw entropy values are very close, we scaled the y‑axis by a factor of 103. We could observe that, across all layers, image‑region entropy for real tokens is lower than hallucinated tokens. However, in some architectures, like shikra, we observe that in certain layers the average entropy is actually higher for real words than for hallucinated terms. This further confirms that the indiscriminate modification of attention heads, as previously discussed, is suboptimal.
@@ D.2 Metric Validation and Comparison
To evaluate the significance of the metric, which independent and non‑normally distributed across real and hallucinated tokens, we apply the Wilcoxon signed‑rank test (Wilcoxon, 1992) as in main text. With $p<0.001$, in Figure 10 we confidently observe in MiniGPT-4 that hallucinated tokens exhibit significantly lower LIAF compared to real tokens.
@@
While we have already confirmed the individual effectiveness of LIAE and LIAF, combining them necessitates conducting a correlation analysis between the two metrics. Since these two metrics do not conform to a normal distribution, we computed the Spearman correlation coefficient between LIAE and LIAF, obtaining $\rho=–0.85$ with $p<0.001$, which indicates that there is a high negative correlation between LIAE and LIAF. To further characterize their relationship, we fitted an isotonic regression—shown as the solid blue curve in Figure 11 left closely follows, yet remains slightly above, the idealized gray dashed line denoting perfect negative correlation. We obtained the same fitting results for attention head localization metrics: Image Attention Entropy (IAE) and Image Attention Focus (IAF) as illustrated in Figure 11 right.
@@
To comprehensively evaluate the capability of the three proposed metrics in detecting abnormal layers, we conducted experiments on the CHAIR dataset using LLaVA, MiniGPT-4, and Shikra, as shown in the table 5. The results demonstrate that selecting LIAE alone as the primary indicator yields the strongest hallucination suppression. However, as LIAF is incorporated, the suppression effect gradually diminishes: once $\alpha$ exceeds 0.5, all three models produce identical detection results. We attribute this to the substantial numerical disparity between LIAE and LIAF, which causes LIAF to increasingly dominate in the mixed metric, whereas LIAE is inherently more sensitive to abnormalities. Therefore, we use LIAE exclusively during detection.
@@
We visualize the distributional differences of IAE when the model generates hallucinated versus factual tokens, as shown in Figure 12. The results reveal that the discrepancies across attention heads are extremely subtle, appearing only beyond the fifth decimal place. We further repeat the CHAIR experiment and find that once IAE is incorporated, the correction process leads the model to malfunction, as it mistakenly identifies and modifies the wrong attention heads. For these reasons, we use IAF exclusively during correction.
@@ D.3 Evaluation of Metric Generalizability
To further verify the validity and effectiveness of our proposed metrics across different model architectures, we conducted supplementary experiments on other model architecture and analyzed the results. We first plot the distributions of LIAE and IAF in Shikra like the main context in Figure 13.
@@
Additionally, we plotted the LIAF curve and distributions in Figure 14 and performed Wilcoxon signed‑rank tests on both metrics for hallucinated versus ground‑truth tokens. At p << 0.001, hallucinated tokens exhibit significantly lower LIAF than ground‑truth tokens.
@@
Furthermore, we evaluated the correlation between LIAE and LIAF and fitted it with isotonic regression, as shown in Figure 15 left. The gray dashed line denotes a perfect negative correlation, and the blue curve represents the fitted regression; we observe a correlation of –0.88 at p<<0.001.
After confirming that LIAE and LIAF generalize well for identifying anomalous layers, we further evaluated the generalizability of Image Attention Entropy (IAE) and Image Attention Focus (IAF) for pinpointing anomalous attention heads. The distribution and significance test of IAE is presented in Figure 12: for each attention head, Image Attention Entropy (IAE) in ground‑truth tokens is significantly lower than in hallucinated tokens (p << 0.001).
Similarly, we applied isotonic regression fitting; as shown in Figure 15 right, although our curve does not perfectly align with the gray dashed line denoting a perfect negative correlation, it still exhibits a clear negative trend, with a correlation coefficient of –0.68 (p<<0.001).
@@ D.4 Visual Processing Layer
In the main text, we conducted an ablation study on the Layer Prior and found that even without incorporating it, our algorithm already achieves state-of-the-art performance. However, introducing the layer prior further reduces hallucination rates. In this section, we provide a more fine-grained analysis of the Layer Prior and introduce the concept of visual processing layers.
We first evaluated the effect of varying the number of corrected layers on three models: LLaVA, MiniGPT-4, and Shikra, measuring changes in hallucination rate and F1 score, as shown in Table 6. The results demonstrate that regardless of the partitioning strategy, our method consistently achieves a substantial reduction in hallucination rates. Moreover, by tuning the number of corrected layers, our approach is able to maintain high F1 scores while suppressing hallucinations, further confirming both the robustness and strong transferability of our algorithm.
@@
We tracked the trajectories of the top $90\%$ percentile logits across all layers of MiniGPT-4 and Shikra (Figure 16 and 17). The curves plateau around layer 26, suggesting that content integration and reasoning are effectively completed by the end of the first 25 layers. Accordingly, our hallucination detection and correction mechanisms are concentrated on these initial layers, which confirm the prior hypothesis.
@@ D.5 Full Procedure of D-LEAF
Algprithm 1 presents the complete D‑LEAF workflow, from layer localization and attention‑head identification to hybrid correction.
@@
17:    for $h\in H$ do
18:     $\text{ISA}_{h}^{(l)}=\text{ISF}_{h}^{(l)}$
19:    end for
20:    $A_{set}=\text{sort}(A_{tot},\text{key}=\text{ISA},\text{asc}=\text{True})$
21:    for $A_{h}\in A_{set}[0:n]$ do
22:     $A_{h,v}=A_{h}[I_{start}:I_{end}]$
23:     $A_{best,v}=A_{set}[-1][I_{start}:I_{end}]$
24:     $A_{h,v}=\gamma A_{best,v}+(1-\gamma)A_{h,v}$
25:    end for
26:   end if
27:  end if
28:  $X=\text{FFN}(A_{tot}V)$
29:  $l=l+1$
30:end while
31:return $X$
Appendix E Case Study
@@
Additional case studies across diverse MLLMs are provided to demonstrate the effectiveness of D-LEAF. Specifically, Figures 1819 and 20 respectively present further instances of hallucination corrections by our method for LLaVA-1.5, MiniGPT-4 and Shikra.

### v2

@@ C.2 LogitLens
C.3 Wilcoxon Signed-rank Test
@@ C.4 Isotonic Regression
C.5 Spearman Correlation Coefficient
@@ D.1 Empirical Analysis on Entropy and Focus
We randomly sampled 500 images from the COCO2014 validation set and, for each image, extracted the model’s attention matrices when generating hallucinated versus ground‑truth tokens, respectively. We then computed the attention scores over the image region for both cases. The results are shown in Fig. 14. We observe that, in both models, across all layers, image‑region attention for ground‑truth tokens is higher than hallucinated tokens.
@@
Similarly, after applying softmax normalization to the attention matrices for ground‑truth and hallucinated tokens, we computed the mean image‑region entropy across all attention heads at each layer and plotted their per‑layer difference as a bar chart in Fig. 15. Because the raw entropy values are very close, we scaled the y‑axis by a factor of 103. We could observe that, across all layers, image‑region entropy for real tokens is lower than hallucinated tokens. However, in some architectures, like shikra, we observe that in certain layers the average entropy is actually higher for real words than for hallucinated terms. This further confirms that the indiscriminate modification of attention heads, as previously discussed, is suboptimal.
@@ D.2 Metric Validation and Comparison
To evaluate the significance of the metric, which independent and non‑normally distributed across real and hallucinated tokens, we apply the Wilcoxon signed‑rank test [wilcoxon1992individual] as in main text. With $p<0.001$, in Fig. 16 we confidently observe in MiniGPT-4 that hallucinated tokens exhibit significantly lower LIAF compared to real tokens.
@@
While we have already confirmed the individual effectiveness of LIAE and LIAF, combining them necessitates conducting a correlation analysis between the two metrics. Since these two metrics do not conform to a normal distribution, we computed the Spearman correlation coefficient between LIAE and LIAF, obtaining $\rho=–0.85$ with $p<0.001$, which indicates that there is a high negative correlation between LIAE and LIAF. To further characterize their relationship, we fitted an isotonic regression—shown as the solid blue curve in Fig. 17 left closely follows, yet remains slightly above, the idealized gray dashed line denoting perfect negative correlation. We obtained the same fitting results for attention head localization metrics: Image Attention Entropy (IAE) and Image Attention Focus (IAF) as illustrated in Fig. 17 right.
@@
To comprehensively evaluate the capability of the three proposed metrics in detecting abnormal layers, we conducted experiments on the CHAIR dataset using LLaVA, MiniGPT-4, and Shikra, as shown in the Tab. 4. The results demonstrate that selecting LIAE alone as the primary indicator yields the strongest hallucination suppression. However, as LIAF is incorporated, the suppression effect gradually diminishes: once $\alpha$ exceeds 0.5, all three models produce identical detection results. We attribute this to the substantial numerical disparity between LIAE and LIAF, which causes LIAF to increasingly dominate in the mixed metric, whereas LIAE is inherently more sensitive to abnormalities. Therefore, we use LIAE exclusively during detection.
@@
We visualize the distributional differences of IAE when the model generates hallucinated versus factual tokens, as shown in Fig. 18. The results reveal that the discrepancies across attention heads are extremely subtle, appearing only beyond the fifth decimal place. We further repeat the CHAIR experiment and find that once IAE is incorporated, the correction process leads the model to malfunction, as it mistakenly identifies and modifies the wrong attention heads. For these reasons, we use IAF exclusively during correction.
@@ D.3 Evaluation of Metric Generalizability
To further verify the validity and effectiveness of our proposed metrics across different model architectures, we conducted supplementary experiments on other model architecture and analyzed the results. We first plot the distributions of LIAE and IAF in Shikra like the main context in Fig. 19.
@@
Additionally, we plotted the LIAF curve and distributions in Fig. 20 and performed Wilcoxon signed‑rank tests on both metrics for hallucinated versus ground‑truth tokens. At p << 0.001, hallucinated tokens exhibit significantly lower LIAF than ground‑truth tokens.
@@
Furthermore, we evaluated the correlation between LIAE and LIAF and fitted it with isotonic regression, as shown in Fig. 21 left. The gray dashed line denotes a perfect negative correlation, and the blue curve represents the fitted regression; we observe a correlation of –0.88 at p<<0.001.
After confirming that LIAE and LIAF generalize well for identifying anomalous layers, we further evaluated the generalizability of Image Attention Entropy (IAE) and Image Attention Focus (IAF) for pinpointing anomalous attention heads. The distribution and significance test of IAE is presented in Fig. 18: for each attention head, Image Attention Entropy (IAE) in ground‑truth tokens is significantly lower than in hallucinated tokens (p << 0.001).
Similarly, we applied isotonic regression fitting; as shown in Fig. 21 right, although our curve does not perfectly align with the gray dashed line denoting a perfect negative correlation, it still exhibits a clear negative trend, with a correlation coefficient of –0.68 (p<<0.001).
@@ D.4 Visual Processing Layer
In the Fig. 12, we conducted an ablation study on the Layer Prior and found that even without incorporating it, our algorithm already achieves state-of-the-art performance. However, introducing the layer prior further reduces hallucination rates. In this section, we provide a more fine-grained analysis of the Layer Prior and introduce the concept of visual processing layers.
We first evaluated the effect of varying the number of corrected layers on three models: LLaVA, MiniGPT-4, and Shikra, measuring changes in hallucination rate and F1 score, as shown in Tab. 5. The results demonstrate that regardless of the partitioning strategy, our method consistently achieves a substantial reduction in hallucination rates. Moreover, by tuning the number of corrected layers, our approach is able to maintain high F1 scores while suppressing hallucinations, further confirming both the robustness and strong transferability of our algorithm.
@@
We tracked the trajectories of the top $90\%$ percentile logits across all layers of MiniGPT-4 and Shikra (Fig. 22 and Fig. 23). The curves plateau around layer 26, suggesting that content integration and reasoning are effectively completed by the end of the first 25 layers. Accordingly, our hallucination detection and correction mechanisms are concentrated on these initial layers, which confirm the prior hypothesis.
@@ D.5 Full Procedure of D-LEAF
Algorithm 1 and Fig. 24 shows the full procedure of D-LEAF.
Figure 24: Architecture of our D-LEAF.
@@
17:    $A_{set}=\text{sort}(A_{tot},\text{key}=\text{IAF},\text{asc}=\text{True})$
18:    for $A_{h}\in A_{set}[0:n]$ do
19:     $A_{h,v}=A_{h}[I_{start}:I_{end}]$
20:     $A_{best,v}=A_{set}[-1][I_{start}:I_{end}]$
21:     $A_{h,v}=\gamma A_{best,v}+(1-\gamma)A_{h,v}$
22:    end for
23:   end if
24:  end if
25:  $X=\text{FFN}(A_{tot}V)$
26:  $l=l+1$
27:end while
28:return $X$
Appendix E Connection with DPO
In this section, we present the conceptual connection between D-LEAF and DPO. Our analysis is based on the theoretical framework of [uppaal2024model, yang2025nullu] and where a simple logistic model is adopted to represent the probability of the output token given the continuing prompt. Similar to the main text, the subsequent analysis will be conducted separately for each layer.
Given an input prompt along with its hallucinated and non-hallucinated responses, the corresponding embeddings are denoted as $x$, $^{+}, yy^{-}$. The loss function optimized in DPO is defined as:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})$ $=-\,\mathbb{E}_{(x,y^{+},y^{-})\sim D}\bigg[\log\sigma\!\left(\beta\log\frac{\pi_{\theta}(y^{+}|x)}{\pi_{\text{ref}}(y^{+}|x)}\right)$ (19)
$\quad-\,\beta\log\frac{\pi_{\theta}(y^{-}|x)}{\pi_{\text{ref}}(y^{-}|x)}\bigg]$
where, $\pi_{\text{ref}}$ corresponds to the reference (or base) probability model generating output $y$ given $x$, $\pi_{\theta}$ is the new probability model (parametrized by $\theta$), $\sigma$ is the logistic function with $\sigma(z)=(1+\exp(-z))^{-1}$, and $\beta>0$ is a hyperparameter. The gradient of the loss $L_{\text{DPO}}$ with respect to $\theta$ at initialization $\pi_{\theta}=\pi_{\text{ref}}$ equals
$\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})\big|_{\pi_{\theta}=\pi_{\text{ref}}}=$ (20)
$-\,\beta\,\mathbb{E}_{(x,y^{+},y^{-})\sim D}\big[\nabla_{\theta}\log\pi$ ${}_{\theta}(y^{+}|x)-\nabla_{\theta}\log\pi_{\theta}(y^{-}|x)\big]\big|_{\pi_{\theta}=\pi_{\text{ref}}}$
For the language decoder of the MLLMs, let $\mathcal{V}$ represent the vocabulary. We begin with a prompt $x\in\mathcal{V}$ and sequentially generate M next-token predictions $y_{1},y_{2},...,y_{M}\in\mathcal{V}$. At each step, the model predicts the token $y_{m}$based on the input $x_{m}=(x,y_{1},y_{2},...,y_{m-1})$, where $\boldsymbol{x}_{m}$ denotes the encoded representation of the current prompt. We assume a logistic model that generates each continuation $y_{m}$ conditioned on $x_{m}$ defined as follows.
$\pi_{\theta}(y_{m}\mid\boldsymbol{x}_{m})\equiv\pi_{\boldsymbol{W}}(y_{m}\mid\boldsymbol{x}_{m})=Z_{m,\boldsymbol{W}}^{-1}\exp\!\left(\boldsymbol{o}_{y_{m}}^{\top}\boldsymbol{W}\boldsymbol{x}_{m}\right)$ (21)
Here, $\boldsymbol{o}_{y_{m}}$ is the classification vector which we use to get the final word prediction, $\boldsymbol{W}$ is a weight matrix and $Z_{m,\boldsymbol{W}}$ is the normalizing constant:
$Z_{m,\boldsymbol{W}}=\sum_{y\in\mathcal{V}}\exp\!\left(\boldsymbol{o}_{y_{m}}^{\top}\boldsymbol{W}\boldsymbol{x}_{m}\right)$ (22)
In Eq. 21, we assume the classifier is performed with linear transformed module rather than the more common non-linear transformations in the transformer architecture. Based on this model, we then obtain $y=(y_{1},y_{2},...,y_{M})$ from the prompt $x$ as follows:
$\pi_{\theta}(y\mid x)$ $\equiv\pi_{\boldsymbol{W}}(y\mid x)=\prod_{m=1}^{M}\pi_{\boldsymbol{W}}(y_{m}\mid\boldsymbol{x}_{m})$ (23)
$=Z_{\boldsymbol{W}}^{-1}\exp\!\left(\sum_{m=1}^{M}\boldsymbol{o}_{y_{m}}^{\top}\boldsymbol{W}\boldsymbol{x}_{m}\right),$
where $Z_{\boldsymbol{W}}=\prod_{m=1}^{M}Z_{m,\boldsymbol{W}}$. We denote by $\boldsymbol{x}_{m}^{\pm}$, $\boldsymbol{x}_{m}^{\pm}$ and $\boldsymbol{o}_{y_{m}}^{\pm}$ the positive/negative inputs, the corresponding embedding and classification vector for the positive/negative continuation respectively. Plugging this into Eq. 20, the first step DPO update has gradient:
$\nabla_{\boldsymbol{W}}\mathcal{L}_{\text{DPO}}(\pi_{\boldsymbol{W}};\pi_{\text{ref}})\big|_{\pi_{\boldsymbol{W}}=\pi_{\text{ref}}}$ (24)
$=-\,\beta\,\mathbb{E}_{(x,y^{+},y^{-})\sim D}\Bigg[\sum_{m=1}^{M}\Big(\boldsymbol{o}$ ${}_{y_{m}^{+}}(\boldsymbol{x}_{m}^{+})^{\top}-\boldsymbol{o}_{y_{m}^{-}}(\boldsymbol{x}_{m}^{-})^{\top}\Big)\Bigg].$
Note that the normalization factors $Z_{m,\boldsymbol{W}}$ (and hence $Z_{\boldsymbol{W}}$) are omitted when we take the difference of the gradients of the log-probabilities. With $N$ pairs of inputs in $\mathcal{D}$, and we consider the case $M=1$, the DPO gradient will be an average over all the pairs:
$\nabla_{\boldsymbol{W}}\mathcal{L}_{\text{DPO}}(\pi_{\boldsymbol{W}};\pi_{\text{ref}})\big|_{\pi_{\boldsymbol{W}}=\pi_{\text{ref}}}$ (25)
$=-\frac{\beta}{N}\sum_{i=1}^{N}\Big($ $\boldsymbol{o}_{y_{i}^{+}}(\boldsymbol{x}_{i}^{+})^{\top}-\boldsymbol{o}_{y_{i}^{-}}(\boldsymbol{x}_{i}^{-})^{\top}\Big),$
where the extra index $i$ indicates the $i$-th sample pair. The Eq. 25 corresponds to the Theorem 1 in our main paper, which is
$\nabla_{\boldsymbol{W}}\mathcal{L}_{\text{DPO}}$ $=-\frac{\beta}{N}\sum_{i=1}^{N}\Big(\boldsymbol{o}_{y_{i}^{+}}(\boldsymbol{x}_{i}^{+})^{\top}-\boldsymbol{o}_{y_{i}^{-}}(\boldsymbol{x}_{i}^{-})^{\top}\Big)$ (26)
$=-\frac{\beta}{N}$ $\sum_{i=1}^{N}\Big(\underbrace{\boldsymbol{o}_{y_{i}}(\boldsymbol{x}_{i}^{+}-\boldsymbol{x}_{i}^{-})^{\top}}_{\text{feature difference}}+\underbrace{(\boldsymbol{o}_{y_{i}^{+}}-\boldsymbol{o}_{y_{i}^{-}})(\boldsymbol{x}_{i}^{-})^{\top}}_{\text{output difference}}\Big).$
Theorem 1 shows that D-LEAF improves the embeddings by adjusting the underperforming attention heads, making them closer to the embeddings produced when the model generates the preferred output. Therefore, both D-LEAF and DPO mitigate hallucinations by modifying the embeddings so that the model’s outputs align more closely with the preferred responses.
Appendix F Case Study
@@
Additional case studies across diverse MLLMs are provided to demonstrate the effectiveness of D-LEAF. Specifically, Fig. 25, Fig. 26 and Fig. 27 respectively present further instances of hallucination corrections by our method for LLaVA-1.5, MiniGPT-4 and Shikra.